Today we have performed few tests that could be useful for the future commissioning of O5.
1) We opened the RFC loop and we checkd the locking trigger and the error signal. Figure 1 shows the relevant signals for the lock in reflection. The error signals show clear ringing, the usual trigger is too noisy to be used, and the previous trigger is too slow for the purpose (it was low-passed because otherwise there was too much noisy). Conclusion: we can't lock without the RFC locked.
2) We have checked that when we lock the green we see the beam in the cameras. Then we have also misaligned the NE mirror and we have moved the NI and checked if we could see the markers on the border of the mirror. Figures 2 and 3 show the marker with and withuot the beam. Notice that the signal is dominated by the Photon calibrator probably the contrast will improve if we turn it off.
3)Then we locked using only the green (without the IR), splitting the correction of the loop between the input mirror and the VCO. We had already tested the reallocation, but this time we tried with the control loop that we use with the beating, since it improves the accuracy. Figures 4 and 5 show the arm transmission with both filters, it can clearly be seen that with the second we can at least stay in the resonance of the arms in a stable way. Figure 6 shows the spectrum of the loop in both cases. The second filter uses a gain of 200.